Tia saving of Rp 2,000,000 on January 1, 2010 at a bank which provides a simple interest of 9% p.a. If she expects to have savings amounting to Rp 4,000,000 for travel and shopping to Singapore in the new year January 1, 2011, how much savings she had deposited again on July 1, 2010?
1
Expert's answer
2014-10-10T11:54:22-0400
A time deposit is a money deposit at a banking institution that cannot be withdrawn for a certain "term" or period of time (unless a penalty is paid).
The formula of simple interest is:
FV = PV*(1 + i*t), where PV - initial sum, i - interest rate, t - number of periods.
In our case:
4,000,000 = 2,000,000*(1 + 0.09*1) + x*(1 + 0.09*0.5)
4,000,000 = 2,000,000*1.09 + x*1.045
x = 1,820,000/1.045 = $1,741,626.79 should be deposited again on July 1, 2010.